# Stochastic Multi-round Submodular Optimization with Budget

### v1

@@ 4 Semi-Adaptive Approximation Algorithms
Despite our semi-adaptive policy decides the budget allocation in advance, it strongly takes into account the heterogeneity of the input instance over all rounds. Indeed, if the budget was allocated without considering the structure of the input instance (e.g., by distributing uniformly the budget), the resulting approximation could be non-constant.
@@
To show this, assume that the budget allocation is determined by a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ that depends on the number of rounds $T$ and the budget $B$ only, and let $t^{*}$ be the index $t$ minimizing $b_{t}$ (among $b_{1},b_{2},\ldots,b_{T}$). Now, consider an instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ where $|V|=B=T$ and $|H|=1$. As $|H|=1$, the underlying problem is deterministic, thus we can avoid at all the presence of states. Each $f_{t}$ is defined as follows: $f_{t}(S)=|S|$ for any $S\subseteq V$ if $t=t^{*}$ (that is, the state function at round $t^{*}$ counts the number of selected items) and $f_{t}(S)=0$ if $t\neq t^{*}$ (that is, the state function at any round $t\neq t^{*}$ is zeroed).
@@
We have that the optimal policy $\pi^{*}$ for $I$ simply selects all items at round $t^{*}$ (that is, the unique round that produces a positive value of the aggregated state function) and does not select any item at rounds $t\neq t^{*}$; the resulting value of $\pi^{*}$ is $\sigma(\pi^{*})=B=T$. Instead, the value guaranteed by any policy $\pi$ that allocates the budget according to vector $\vec{b}$ is $\sigma(\pi)=b_{t^{*}}=\min_{t\in[T]}b_{t}\leq B/T=T/T=1$. Thus, policy $\pi$ is only $\alpha$-approximate with $\alpha=b_{t^{*}}/B\leq 1/T$ (that is non-constant due to dependence on $T$).
4.1 Estimating State Functions
@@
In order to solve the desired maximization problems, we first need to estimate the expected value of the state functions achieved when selecting an item. However, this is not an easy task due to the stochastic nature of the problem. We first introduce a standard Monte-Carlo algorithm, called ${\sf Oracle1}$, that can be used to estimate the expected value of the state function when selecting an item, with arbitrary precision and arbitrarily high probability: such algorithm can be used to find an item that approximately maximizes the expected influence spread of a state functions, under a given partial state.
@@
Then, we provide another Monte-Carlo algorithm, called ${\sf Oracle2}$, that can be used to estimate the expected increment of the state function achieved by a given adaptive policy $\pi$ when selecting the $j$-th item (in order of selection) at any round $t$, with arbitrary precision and arbitrarily high probability. By exploiting the Chernoff-Hoeffding’s inequality [25], we show that, for any $\delta,\xi>0$ both algorithms can be set to guarantee a precision with an additive error of at most $\delta$ with probability at least $1-\xi$, and to run in polynomial time w.r.t. to the size of the input instance, $1/\delta$ and $1/\xi$. See the SM for a complete description of algorithms ${\sf Oracle1}$, ${\sf Oracle2}$ and their properties.
@@ 4.1.1 Oracle1
${\sf Oracle1}$ takes in input an integer $q\geq 1$, an instance $I$ of the SBMSm problem, a round $t\in[T]$, a partial state $\langle S_{t}\rangle$ for round $t$, an item $v$, and picks $q$ samples $X_{t,v,\langle S\rangle}^{1},\ldots,X_{t,v,\langle S\rangle}^{q}$ of the random variable $\bm{X}_{t,v,\langle S\rangle}:=f_{t}(S_{t}\cup\{v\},\bm{\eta}_{t})-f_{t}(S_{t},\bm{\eta}_{t})$ conditioned by the observation of $\langle S_{t}\rangle$. Then, ${\sf Oracle1}$ returns the value ${\sf Oracle1}(q,I,t,\langle S_{t}\rangle,v):=\tilde{\Delta}_{t}(v|\langle S_{t}\rangle):=\sum_{j=1}^{q}X_{t,v,\langle S\rangle}^{j}/q$.
@@
Let $t\in[T]$, $\langle S\rangle$ be a partial state at round $t$ and, for any $v\in V$, let $\bm{X}_{t,v,\langle S\rangle}$ and $X^{1}_{t,v,\langle S\rangle},\ldots,X^{q}_{t,v,\langle S\rangle}$ be the random quantities considered in the definition of Oracle1. As $\bm{X}_{t,v,\langle S\rangle}$ has values in interval $[0,\Lambda]$ (by the polynomial boundedness of the state function), we have that
@@
$\geq\mathbb{P}\left[\Bigg{|}\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]-\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}\Bigg{|}\leq\delta/2\ \forall v\in V\right]$
@@
where (6) holds as $-\tilde{\Delta}(v^{*}|\langle S_{t}\rangle)+\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)\geq 0$ (by the optimality of $\tilde{v}$ for $\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)$), and (7) follows from (5), and shows property (i).
As for property (ii) we first observe that each oracle sample $X^{j}_{t,v,\langle S_{t}\rangle}$ can be computed in polynomial time. Indeed, for a given partial state $\langle S_{t}\rangle$, we can flip a biased coin with probability $\mu(e)$ for any atom $e\notin R(\langle S_{t}\rangle)$ (i.e., any atom not yet explored by the reachability function $R$), to obtain a state $\eta_{t}\sim\bm{\eta}_{t}$ conditioned by the observation of $\langle S_{t}\rangle$. Then, one can compute $f(S_{t}\cup\{v\},\eta_{t})$ and $f(S_{t},\eta_{t})$ in polynomial time (by exploiting the polynomial time computability of $R$ and $f$) and set $X^{j}_{t,v,\langle S_{t}\rangle}:=f(S_{t}\cup\{v\},\eta_{t})-f(S_{t},\eta)$. Furthermore, $q(\delta,\xi)$ is polynomial, since $\Lambda$ is polynomial in the size of the input. Then $q$ is also polynomial if $q\leq P(q(\delta,\xi))$. ∎
@@ 4.2 Single-Round Adaptive Policy
3:     $v\leftarrow\arg\max_{v^{\prime}\in V}{\sf Oracle1}(q(\delta,\xi),I,t,\langle S_{t}\rangle,v^{\prime})$;$\Delta_{t}(v|\langle S_{t},\eta_{t}\rangle)\geq\max_{v^{\prime}\in V\setminus S_{t}}\Delta_{t}(v^{\prime}|\langle S_{t},\eta_{t}\rangle)-\delta$ with prob. at least $1-\xi$ by Proposition 2 (see Section 4.1.1)
@@
Notice that ${\sf SingleGr}_{\delta,\xi}$ is equivalent to an approximate variant of the adaptive greedy policy defined by [22], that provides a $\left(1-\frac{1}{e}\right)$-approximation to the best policy when considering a unique round.
@@ 4.3 Budget Allocation Algorithm
In order to allocate the budget over rounds, we consider an algorithm called ${\sf BudgetGr}_{\delta,\epsilon}(I)$ (see Algorithm 2 for the pseudo-code), that greedily assigns (in a non-adaptive way) each unit of budget to the round $t$ that maximizes the expected increment of the state function achieved by the greedy single-round (adaptive) policy; the above expected increment is measured with arbitrarily precision and arbitrarily high probability by algorithm Oracle2, in which the single-greedy policy is executed many times to collect a sufficiently high number of samples on the state function increments.
@@
7:     $\overline{\Delta}_{t,n+1}\leftarrow-\infty$;
@@
9:  % By Proposition 3 (see Section 4.1.2) applied to ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, “$|\Delta_{t,i}-\tilde{\Delta}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$” holds with prob. at least $1-\xi$, where $\Delta_{t,i}$ denotes the (exact) expected increment of state function $f_{t}$, achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item, in order of selection. Furthermore, as $\Delta_{t,i}$ is non-increasing in $i$ up to a small addend (by adaptive-submodularity), we also have that the distance $|\Delta_{t,i}-\overline{\Delta}_{t,i}|$ is small with prob. at least $1-\xi$.
@@ 4.4 Multi-Round Adaptive Policy
Let $\vec{b}:=(b_{1},\ldots,b_{T})$ be the output of ${\sf BudgetGr}_{\delta,\epsilon}(I)$. The greedy multi-round (adaptive) policy ${\sf MultiGr}_{\delta,\xi}(I)$ (see Algorithm 1) applies the greedy single-round policy with budget $b_{t}$ at each round $t\in[T]$ to determine a subset $S_{t}$, and returns the aggregated subset $\vec{S}=(S_{1},\ldots,S_{T})$.
@@
Theorem 1 shows that ${\sf MultiGr}_{\delta,\xi}(I)$ is a $(1/2,\epsilon)$-approximate policy for any $\epsilon>0$, for opportune choices of $\delta$ and $\xi$.
@@
Let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},R,(f_{t})_{t\in[T]})$ be an input instance of SBMSm. The greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ is a $\left(1/2-\epsilon\right)$-approximate policy for $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ and any constant $c$ such that $40<c\leq 1/$ 4, and it is polynomial in the size of $I$ and $1/\epsilon$.
@@
Before to provide the full proof of this theorem (provided in the next section), we would like to describe the main proof arguments necessary to get a $1/2$-approximation assuming that the probabilistic oracles are exact (i.e., assuming that Oracle1 and Oracle2 can be run with $\delta:=\xi:=0$).
Consider the following artificial adaptive policy $\hat{\pi}^{*}$ that works in two phases: in the first phase $\hat{\pi}^{*}$ selects $B$ items by running the greedy multi-round policy ${\sf MultiGr}_{0,0}(I)$; in the second phase, $\hat{\pi}^{*}$ restarts from round $1$, and by observing the partial state of the previous item selections, it selects further $\min\{nT-B,B\}$ items in an optimal adaptive way over $T$ rounds. By construction, the performance of $\hat{\pi}^{*}$ is not worse than that of an optimal adaptive policy $\pi^{*}$ selecting $B$ items only. By exploiting the adaptive submodularity, one can show that the expected increment of the state functions achieved by any item selected in the first phase is at least equal as that achieved by each item selected in the second phase. Thus, considering that at most $B$ items are chosen in the second phase, and exactly $B$ items are chosen in the first one, we necessarily have that the expected increment of the aggregated state function achieved in the first phase (that is, $\sigma({\sf MultiGr}_{0,0}(I))$) is at least equal as that achieved in the second phase (that is, $\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}_{0,0}(I))$). Then, by rearranging the previous inequality, we get $\sigma({\sf MultiGr}_{0,0}(I))\geq\frac{1}{2}\sigma(\hat{\pi}^{*})\geq\frac{1}{2}\sigma(\hat{\pi})$, and this shows the desired approximation.

### v2

@@ 4 Partially Adaptive Approximation Algorithm
Despite a partially adaptive policy decides the budget allocation in advance, it must strongly takes into account the heterogeneity of the input instance over all rounds, in order to obtain a good approximation of the optimal adaptive value. Indeed, if the budget was allocated without considering the structure of the input instance (e.g., by distributing uniformly the budget), the resulting approximation could be non-constant.
@@
To show this, assume that the budget allocation is determined by a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ that depends on the number of rounds $T$ and the budget $B$ only, and let $t^{*}$ be the index $t$ minimizing $b_{t}$ (among $b_{1},b_{2},\ldots,b_{T}$). Now, consider an instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ where $|V|=B=T$ and $|H|=1$. As $|H|=1$, the underlying problem is deterministic, thus we can avoid at all the presence of states. Each $f_{t}$ is defined as follows: $f_{t}(S)=|S|$ for any $S\subseteq V$ if $t=t^{*}$ (that is, the objective function at round $t^{*}$ counts the number of selected items) and $f_{t}(S)=0$ if $t\neq t^{*}$ (that is, the objective function at any round $t\neq t^{*}$ is zeroed).
@@
We have that the optimal policy $\pi^{*}$ for $I$ simply selects all items at round $t^{*}$ (that is, the unique round that produces a positive value of the aggregated objective function) and does not select any item at rounds $t\neq t^{*}$; the resulting value of $\pi^{*}$ is $\sigma(\pi^{*})=B=T$. Instead, the value guaranteed by any policy $\pi$ that allocates the budget according to vector $\vec{b}$ is $\sigma(\pi)=b_{t^{*}}=\min_{t\in[T]}b_{t}\leq B/T=T/T=1$. Thus, policy $\pi$ is only $\alpha$-approximate with $\alpha=b_{t^{*}}/B\leq 1/T$ (that is non-constant due to dependence on $T$).
Below, we provide the detailed description of such policy and the analysis of its approximation factor.
4.1 Estimating Objective Functions
@@
In order to solve the desired maximization problems through our greedy partially adaptive policy, we first need to estimate the expected value of the objective functions achieved when selecting an item. However, this is not an easy task due to the stochastic nature of the problem. We first introduce a standard Monte-Carlo algorithm, called ${\sf Oracle1}$, that can be used to give a good estimation, with high probability, of the expected value of the objective function when selecting an item: such algorithm can be used to find an item that approximately maximizes the expected influence spread of an objective function, under a given partial state.
@@
Then, we provide another Monte-Carlo algorithm, called ${\sf Oracle2}$, that can be used to estimate the expected increment of the objective function achieved by a given adaptive policy $\pi$ when selecting the $j$-th item (in order of selection) at any round $t$; as in Oracle1, the precision of such estimation will be good with high probability. By exploiting the Chernoff-Hoeffding’s inequality [31], we show that, for any $\delta,\xi>0$ both algorithms can be set to guarantee a precision with an additive error of at most $\delta$ with probability at least $1-\xi$, and to run in polynomial time w.r.t. to the size of the input instance, $1/\delta$ and $1/\xi$. See the SM for a complete description of algorithms ${\sf Oracle1}$, ${\sf Oracle2}$ and their properties.
@@ 4.1.1 Oracle1
${\sf Oracle1}$ takes in input an integer $q\geq 1$, an instance $I$ of the SBMSm problem, a round $t\in[T]$, a partial state $\langle S_{t}\rangle$ for round $t$, an item $v$, and picks $q$ samples $X_{t,v,\langle S_{t}\rangle}^{1},\ldots,X_{t,v,\langle S_{t}\rangle}^{q}$ of the random variable $\bm{X}_{t,v,\langle S\rangle}:=f_{t}(S_{t}\cup\{v\},\bm{\eta}_{t})-f_{t}(S_{t},\bm{\eta}_{t})$ conditioned by the observation of $\langle S_{t}\rangle$. Then, ${\sf Oracle1}$ returns the value ${\sf Oracle1}(q,I,t,\langle S_{t}\rangle,v):=\tilde{\Delta}_{t}(v|\langle S_{t}\rangle):=\sum_{j=1}^{q}X_{t,v,\langle S_{t}\rangle}^{j}/q$.
@@
Let $t\in[T]$, $\langle S\rangle$ be a partial state at round $t$ and, for any $v\in V$, let $\bm{X}_{t,v,\langle S\rangle}$ and $X^{1}_{t,v,\langle S\rangle},\ldots,X^{q}_{t,v,\langle S\rangle}$ be the random quantities considered in the definition of Oracle1. As $\bm{X}_{t,v,\langle S\rangle}$ has values in interval $[0,\Lambda]$ (by the polynomial boundedness of the objective function), we have that
@@
$=\mathbb{P}\left[\left|\mathbb{E}\left[\bm{X}_{t,v,\langle S_{t}\rangle}\right]-\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}\right|\leq\delta/2\ \forall v\in\{\tilde{v},v^{*}\}\right]$ (7)
$\geq\mathbb{P}\left[\Bigg{|}\mathbb{E}\left[\bm{X}_{t,v,\langle S_{t}\rangle}\right]-\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S_{t}\rangle}}{q}\Bigg{|}\leq\delta/2\ \forall v\in V\right]$
@@
where (6) holds as $-\tilde{\Delta}(v^{*}|\langle S_{t}\rangle)+\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)\geq 0$ (by the optimality of $\tilde{v}$ for $\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)$), (7) holds since $\Delta(v|\langle S_{t}\rangle)=\mathbb{E}\left[\bm{X}_{t,v,\langle S_{t}\rangle}\right]$ and $\tilde{\Delta}(v|\langle S_{t}\rangle)=\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}$, and (8) follows from (5). Finally, Property (i) follows from (5).
Property (ii) immediately follows from: (a) the polynomial boundedness property, that polynomially relates the quantity $\Lambda$ (appearing in $q(\delta,\xi)$) to the size of the input instance; (b) the existence of an oracle that picks samples drawn from the underlying probability distributions (whose existence has been implicitly assumed in the model definition of Section 2). ∎
@@ 4.2 Single-Round Adaptive Policy
3:     $v\leftarrow\arg\max_{v^{\prime}\in V}{\sf Oracle1}(q(\delta,\xi),I,t,\langle S_{t}\rangle,v^{\prime})$;$\Delta_{t}(v|\langle S_{t},\eta_{t}\rangle)\geq\max_{v^{\prime}\in V\setminus S_{t}}\Delta_{t}(v^{\prime}|\langle S_{t},\eta_{t}\rangle)-\delta$ with prob. at least $1-\xi$ by Proposition 2.
@@
Notice that ${\sf SingleGr}_{\delta,\xi}$ is equivalent to an approximate variant of the adaptive greedy policy defined by [26], that provides a $\left(1-1/e-\epsilon\right)$-approximation to the best policy when considering a unique round.
@@ 4.3 Budget Allocation Algorithm
In order to allocate the budget over rounds, we consider an algorithm called ${\sf BudgetGr}_{\delta,\epsilon}(I)$ (see Algorithm 2 for the pseudo-code), that greedily assigns (in a non-adaptive way) each unit of budget to the round $t$ that maximizes the expected increment of the objective function achieved by the greedy single-round (adaptive) policy; the above expected increment is well-approximated with high probability by algorithm Oracle2, in which the single-greedy policy is executed many times to collect a sufficiently high number of samples on the objective function increments.
@@
7:     $\overline{\Delta}_{t,n+1}\leftarrow-\infty$; % This is done to impose the unfeasibility of selecting $n+1$ items at each round $t$.
@@
9:  % By Proposition 3 applied to ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, the event “$|\Delta_{t,i}-\tilde{\Delta}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$” holds with prob. at least $1-\xi$, where $\Delta_{t,i}$ denotes the (exact) expected increment of objective function $f_{t}$, achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item, in order of selection. Furthermore, as $\Delta_{t,i}$ is non-increasing in $i$ up to a small addend (by adaptive-submodularity), we also have that the distance $|\Delta_{t,i}-\overline{\Delta}_{t,i}|$ is small with prob. at least $1-\xi$.
@@ 4.4 Multi-Round Adaptive Policy
Let $\vec{b}:=(b_{1},\ldots,b_{T})$ be the output of ${\sf BudgetGr}_{\delta,\epsilon}(I)$. The greedy multi-round (adaptive) policy ${\sf MultiGr}_{\delta,\xi}(I)$ (see Algorithm 3 for the pseudo-code) applies the greedy single-round policy with budget $b_{t}$ at each round $t\in[T]$ to determine a subset $S_{t}$, and returns the aggregated subset $\vec{S}=(S_{1},\ldots,S_{T})$.
@@
Theorem 1 shows that, for any fixed $\epsilon>0$ and opportune choices of $\delta$ and $\xi$, ${\sf MultiGr}_{\delta,\xi}(I)$ achieves a $(1-1/e-\epsilon)$ approximation to the optimal adaptive policy. In the following remark, we first explain why the approaches previously adopted to deal with the single-round case [26] could not be directly applied to the general multi-round framework to obtain the same approximation, thus showing that a new approach like ours was somehow necessary.
Remark 2.
As the adaptive greedy algorithm of [26] returns a $(1-1/e-\epsilon)$-approximation for (adaptive) single-round submodular maximization problems, one could reasonably believe that their approach guarantees the same approximation for the multi-round generalization considered in this work; if this was the case, our partially adaptive algorithm would not guarantee a better approximation than their fully adaptive algorithm, and it would be of interest only because partial adaptivity is easier to implement in various realistic contexts. However, a direct application of the adaptive greedy algorithm of [26] would require to represent the input multi-round instance as a union of $T$ independent single-round sub-instances (i.e., one for each round) and, similarly as in the greedy-single round policy (see Algorithm 1), would greedily and adaptively select $B$ items (in $B$ steps), possibly moving back and forth between rounds. Despite our model allows to move from one round to the subsequent ones, it forbids to return to selecting items in previously abandoned rounds (as this would mean go back in time and cancel the choices previously made). Thus, due to this constraint, the greedy choices suggested by the algorithm of [26] may not be valid when applied to our multi-round framework.
Furthermore, if we apply the algorithm of [26] by restricting the greedy choice, at each step, to items of the current or the subsequent rounds, the resulting approximation could be arbitrarily low (i.e., inefficient). To show this, we consider a simple instance $I$ defined as follows: we have two rounds, a set of $n$ items and a budget $B:=n+1$; a special item $v^{*}$ has weight weight 1 for each round, and the remaining items have weight $1/2$ at round $1$ and weight 00 at round $2$; the objective function, at each round, simply counts the total weight of selected items. We observe that the considered instance can be easily modelled as a (fully deterministic) binary submodular stochastic probing problem (see Section 2.5). An optimal policy for $I$ requires to select all items at round $1$ and the special item at round $2$, and then guarantees a value of $(n-1)/2+2=n/2+3/2$. However, the solution returned by the considered adaptation of the greedy algorithm of [26] selects the special item at round 1, moves to the subsequent round and then selects all the available $n$ items (as it cannot go back to round 1); in such a case the obtained value is $2$, and then only guarantees a $2/(n/2+3/2)=4/(n+3)=O(1/n)$-approximation of the optimal value, that becomes arbitrarily low as $n$ increases.
@@
Let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},R,(f_{t})_{t\in[T]})$ be an input instance of SBMSm. The greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ is a $\left(1-1/e-\epsilon\right)$-approximate policy for $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ and any constant $c$ such that $40<c\leq 1/$ 4, and it is polynomial in the size of $I$ and $1/\epsilon$.
@@
Before to provide the full proof of this theorem (given in the next subsection), we would like to describe the main proof arguments necessary to get a $(1-1/e)$-approximation assuming that the probabilistic oracles are exact (i.e., assuming that Oracle1 and Oracle2 can be run with $\delta:=\xi:=0$).
Sketch of the proof (approximation factor for the case of exact oracles).
For any $k\in\llbracket 0,B\rrbracket$, let $\vec{b}(k):=(b_{1}(k),\ldots,b_{T}(k))$ be the budget vector obtained by algorithm ${\sf BudgetGr}_{\delta,\epsilon}(I)$ (see Algorithm 2) after the assignment of $k$-th unit of budget. Let ${\sf MultiGr}_{0,0}(I)$ be the greedy multi-round policy that uses exact oracles (i.e., with $\delta=\xi=0$), $\pi^{*}$ be an optimal adaptive policy, $\pi(k)$ be the greedy multi-round policy ${\sf MultiGr}_{0,0}(I)$ applied to budget vector $\vec{b}(k)$; we observe that $\pi(B)$ coincides with the policy ${\sf MultiGr}_{0,0}(I)$. Consider the following artificial adaptive policy $\hat{\pi}^{*}(k)$ that works in two phases: in the first phase, $\hat{\pi}^{*}(k)$ runs policy $\pi(k)$; in the second phase, $\hat{\pi}^{*}(k)$ restarts from round $1$ and, by observing the partial state of the previous item selections, selects further $\min\{nT-k,B\}$ items in an optimal adaptive way over all $T$ rounds.
Fix $k\in\llbracket 0,B-1\rrbracket$. By the monotonicity of the objective functions and since $\hat{\pi}^{*}(k)$ selects as many items as in the optimal adaptive policy $\pi^{*}$, the performance of $\hat{\pi}^{*}(k)$ is not worse than that of $\pi^{*}$, that is,
$\sigma(\hat{\pi}^{*}(k))\geq\sigma(\pi^{*}).$ (9)
Furthermore, by exploiting the greedy assignment of the budget chosen by ${\sf BudgetGr}_{\delta,\epsilon}(I)$ and the adaptive submodularity, one can show that $\sigma(\pi(k+1))-\sigma(\pi(k))$ (that is, the expected increment of the objective functions achieved by ${\sf MultiGr}_{0,0}(I)$, when selecting the item associated with the $(k+1)$-th unit of budget assigned by ${\sf BudgetGr}_{\delta,\epsilon}(I)$) is at least equal as the expected increment of the objective functions achieved by each item selected in the second phase of $\hat{\pi}^{*}(k)$. Thus, considering that the overall expected utility achieved in the second phase of $\hat{\pi}^{*}(k)$ is $\sigma(\hat{\pi}^{*}(k))-\sigma(\pi(k))$ and, in such phase, at most $B$ items are selected, we necessarily have that
$\sigma(\pi(k+1))-\sigma(\pi(k))\geq\frac{\sigma(\hat{\pi}^{*}(k))-\sigma(\pi(k))}{B}\geq\frac{\sigma(\pi^{*})-\sigma(\pi(k))}{B},$
where the first inequality can be formally shown by setting $\delta:=\xi:=0$ and $\pi^{\prime}(k):=\pi(k)$ in Lemma 1 (see the full proof of Theorem 1 below) and the last one follows from (9). By rearranging the above inequalities we obtain
$\sigma(\pi(k+1))\geq\left(1-\frac{1}{B}\right)\sigma(\pi(k))+\left(\frac{1}{B}\right)\sigma(\pi^{*}).$ (10)
By applying recursively (10) from $k=0$ to $k=B-1$ we obtain $\sigma(\pi(k))\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}$. Thus, we have
${\sf MultiGr}_{0,0}(I)=\sigma(\pi(B))\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}=\left(\frac{1}{B}\right)\sigma(\pi^{*})\left(\frac{1-\left(1-\frac{1}{B}\right)^{B}}{1-\left(1-\frac{1}{B}\right)}\right)$
$\ \ =\left(1-\left(1-\frac{1}{B}\right)^{B}\right)\sigma(\pi^{*})\geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*}),$ (11)
where (11) follows from the known inequality $(1+x)^{r}<e^{rx}$, holding for any $x\geq-1$ and $r>0$, applied with $x=-1/B$ and $r=B$. By (11), we have that ${\sf MultiGr}_{0,0}(I)$ is a $(1-1/e)$-approximate policy, and this shows the claim. ∎